# Localization phenomena in the random XXZ spin chain

### v1

@@ 3. Basic features of the XXZ spin chain
$\left\lvert{\bf u}-{\bf v}\right\rvert_{1}+d_{H}^{D}\left({\bf v},{\bf w}\right)+\left\lvert{\bf w}-{\bf y}\right\rvert_{1}\geq d_{H}^{D}({\bf x},{\bf y}),$ (4.14)
$\left\lvert{\bf u}-{\bf v}\right\rvert_{1}+d_{H}^{D}\left({\bf v},{\bf w}\right)+\left\lvert{\bf w}-{\bf y}\right\rvert_{1}\geq\left\lvert{\bf u}-{\bf v}\right\rvert_{1}+\frac{1}{2}\left(d_{H}^{D}\left({\bf v},{\bf y}\right)+\left\lvert{\bf w}-{\bf y}\right\rvert_{1}\right).$
@@
$\sum_{{\bf u}\in\mathcal{P}_{N}(D):x\in{\bf u}}\ \sum_{{\bf v},{\bf w}\in\mathcal{P}_{N,\lceil q\rceil}(D)}\mathrm{e}^{-c_{q}\left\lvert{\bf u}-{\bf v}\right\rvert_{1}}\mathrm{e}^{-c_{q}d^{D}_{H}\left({\bf v},{\bf w}\right)}\mathrm{e}^{-c_{q}\left\lvert{\bf w}-{\bf y}\right\rvert_{1}}$ (4.15)
$\qquad\leq e^{-\frac{c_{q}}{2}d_{H}^{D}({\bf x},{\bf y})}\sum_{{\bf u}\in\mathcal{P}_{N}(D):}\ \sum_{{\bf v},{\bf w}\in\mathcal{P}_{N,\lceil q\rceil}(D)}e^{-\frac{c_{q}}{2}\left\lvert{\bf u}-{\bf v}\right\rvert_{1}}e^{-\frac{c_{q}}{4}d_{H}^{D}\left({\bf v},{\bf w}\right)}e^{-\frac{c_{q}}{4}\left\lvert{\bf w}-{\bf y}\right\rvert_{1}}$
@@
using $\mathbf{y}\in\mathcal{P}_{N,\lceil q\rceil}$, (B.10) and (B.3) in the last step.
@@
using a large deviation estimate.
Let $\bar{\mu}=\mathbb{E}\left\{\omega_{0}\right\}$, and assume $N\lambda\bar{\mu}>2\lceil q\rceil\left(1-\tfrac{1}{\Delta}\right)$. Then the standard large deviations estimate (recall (3.2)) gives
@@
where we used $\left\lvert\mathcal{P}_{N,\lceil q\rceil}(S)\right\rvert=\operatorname{tr}Q_{\leq k}^{S}$, (3.11), and $\left\lvert S\right\rvert\leq N(2N+1)$. Moreover, on the complimentary event $\left(\mathcal{E}_{N}^{S}\right)^{c}$ we have
@@
so we can use [EKS1, Proposition 4.1] to obtain the Combes-Thomas bound
$\sup_{z\in\mathds{H}_{\lceil q\rceil}}\left\lvert G^{S}_{z}({\bf x},{\bf y})\right\rvert\leq C_{q}e^{-c_{q}\left\lvert{\bf x}-{\bf y}\right\rvert_{1}}\quad\text{for all}\quad\mathbf{x},\mathbf{y}\in\mathcal{P}_{N,\lceil q\rceil}(S).$ (4.22)
To show (4.17), we start by observing that for $\nu\in\sigma_{q}(H^{D})$ we have
$\pi_{{}_{\nu}}\phi_{\bf x}=\pi_{\nu}\left(H^{S,S^{c}}-\nu\right)R_{\nu}^{S,S^{c}}\phi_{\bf x}=\pi_{\nu}\Gamma^{S}R_{\nu}^{S,S^{c}}\phi_{\bf x}.$ (4.23)
By the construction of $S$ we have
@@
It follows that on the complimentary event $\left(\mathcal{E}_{N}^{S}\right)^{c}$ we have
$\sum_{\nu\in\sigma_{q}(H^{D})}\left\lvert\langle\phi_{\bf y},\pi_{\nu}\phi_{\bf x}\rangle\right\rvert$ $\leq\sum_{\nu\in\sigma_{q}(H^{D})}\sum_{{\bf u}\in\mathcal{P}_{N}^{\partial_{ex}}(D)}\left\lvert\langle\phi_{\bf y},\pi_{\nu}\phi_{\bf u}\rangle\right\rvert\left\lvert\langle\phi_{\bf u},\Gamma^{S}R_{\nu}^{S}\phi_{\bf x}\rangle\right\rvert$ (4.25)
$\leq C_{q}\sum_{\nu\in\sigma_{q}(H^{D})}\sum_{{\bf u}\in\mathcal{P}_{N}^{\partial_{ex}}(D)}e^{-c_{q}\left\lvert{\bf u}-{\bf x}\right\rvert_{1}}\left\lvert\langle\phi_{\bf y},\pi_{\nu}\phi_{\bf u}\rangle\right\rvert,$
@@
$\mathcal{P}_{N}^{\partial_{ex}}(D)=\mathcal{P}_{N}(D,\partial_{ex}^{\Lambda}D)=\left\{{\bf u}\in\mathcal{P}_{N}(D),{\bf u}\cap\partial_{ex}^{D}S\neq\emptyset\right\}.$ (4.26)
@@
$\sum_{\nu\in\sigma_{q}(H^{D})}\left\lvert\langle\phi_{\bf y},\pi_{\nu}\phi_{\bf x}\rangle\right\rvert\leq C_{q}\sum_{{\bf u}\in\Psi_{N}}e^{-c_{q}\left\lvert{\bf u}-{\bf x}\right\rvert_{1}}.$ (4.28)
Since for any ${\bf u}\in\mathcal{P}_{N}^{\partial_{ex}}(D)$ we have $\left\lvert{\bf u}-{\bf x}\right\rvert_{1}\geq N$, we deduce from (4.28) and (B.3) (recall $\mathbf{y}\in\mathcal{P}_{N,\lceil q\rceil}(D)$), that on $\mathcal{E}_{N}^{c}$ we have
@@
$\pi_{\nu}^{D}(\mathbf{x},\mathbf{y})=C_{q}\sum_{\mathbf{u},\mathbf{v}\in\mathcal{P}_{N,\lceil q\rceil}(D)}\widehat{G}_{\lceil q\rceil,\nu}(\mathbf{x},\mathbf{u})\pi_{\nu}^{D}(\mathbf{u},\mathbf{v}){\widehat{G}}_{\lceil q\rceil,\nu}(\mathbf{v},\mathbf{y})$ (4.32)
@@
$\mathcal{Q}^{D}(\mathbf{x},\mathbf{y})$ $\leq\sum_{\mathbf{u},\mathbf{v}\in\mathcal{P}_{N,\lceil q\rceil}(D)}\left\lvert\widehat{G}_{\lceil q\rceil,\nu}(\mathbf{x},\mathbf{u})\right\rvert\mathcal{Q}^{D}(\mathbf{u},\mathbf{v})\left\lvert\widehat{G}_{\lceil q\rceil,\nu}(\mathbf{v},\mathbf{y})\right\rvert$ (4.33)
$\leq C_{q}\sum_{\mathbf{u},\mathbf{v}\in\mathcal{P}_{N,\lceil q\rceil}(D)}\mathrm{e}^{-c_{q}\left\lvert\mathbf{x}-\mathbf{u}\right\rvert_{1}}\mathcal{Q}^{D}(\mathbf{u},\mathbf{v})\mathrm{e}^{-c_{q}\left\lvert\mathbf{v}-\mathbf{y}\right\rvert_{1}},$
@@
$\mathbb{E}_{D}\left\{\mathcal{Q}^{D}(\mathbf{x},\mathbf{y}\right\}$ $\leq C_{q}N^{2\lceil q\rceil}\sum_{\mathbf{u},\mathbf{v}\in\mathcal{P}_{N,\lceil q\rceil}(D)}\mathrm{e}^{-c_{q}\left\lvert\mathbf{x}-\mathbf{u}\right\rvert_{1}}\mathrm{e}^{-c_{q}d_{H}^{D}(\mathbf{u},\mathbf{v})}\mathrm{e}^{-c_{q}\left\lvert\mathbf{v}-\mathbf{y}\right\rvert_{1}}$ (4.34)
$\leq C_{q}\mathrm{e}^{-c^{\prime}_{q}d_{H}^{D}(\mathbf{x},\mathbf{y})}\sum_{\mathbf{u},\mathbf{v}\in\mathcal{P}_{N,\lceil q\rceil}(D)}\mathrm{e}^{-c^{\prime}_{q}\left\lvert\mathbf{x}-\mathbf{u}\right\rvert_{1}}\mathrm{e}^{-c^{\prime}_{q}\left\lvert\mathbf{v}-\mathbf{y}\right\rvert_{1}}$
@@
where we used (B.2) twice. The Lemma is proven. ∎
@@
Let $q\ \in\frac{1}{2}\mathbb{N}$, $1\leq q$, and assume the the induction hypothesis, that is, Theorem 2.4 is proven for $q-\frac{1}{2}$. Let $\Lambda\subset\mathbb{Z}$ and $N\in\mathbb{N}$. Then for all $\mathbf{x},\mathbf{y}\in\mathcal{P}_{N,\lceil q\rceil}$ and any finite connected set $D\subset\Lambda$ satisfying $(\mathbf{x}\cup\mathbf{y})_{D}\neq\emptyset$ and $(\mathbf{x}\cup\mathbf{y})_{D^{c}}\neq\emptyset$ we have
$\sup_{z\in\mathds{H}_{q}}\mathbb{E}_{\Lambda}\left\{\left\lvert G^{D,D^{c}}_{z}({\bf x},{\bf y})\right\rvert^{s}\right\}\leq C_{q}\left\lvert D\right\rvert^{2\lceil q\rceil}\mathrm{e}^{-c_{q-\frac{1}{2}}{d_{H}^{\Lambda}({\bf x},{\bf y})}}\;\ \text{for all}\;\ \mathbf{x},\mathbf{y}\in\mathcal{P}_{N,\lceil q\rceil}(D).$ (4.35)
@@
Let $D\subset\Lambda$ be finite and connected, and let $\mathbf{x},\mathbf{y}\in\mathcal{P}_{N,\lceil q\rceil}(\Lambda)$, with $(\mathbf{x}\cup\mathbf{y})_{D}\neq\emptyset$ and $(\mathbf{x}\cup\mathbf{y})_{D^{c}}\neq\emptyset$. We only need to consider the case $11\leq\left\lvert\mathbf{x}_{D}\right\rvert=\left\lvert\mathbf{y}_{D}\right\rvert\leq N-$ 1, as otherwise $G^{D,D^{c}}_{z}({\bf x},{\bf y})=0$.
@@
Given $z\in\mathds{H}_{q}$, let $E=\operatorname{Re}z$ and $a=\operatorname{Im}z$. Setting $r=d_{H}^{\Lambda}\left({\bf x},{\bf y}\right)-1$, and taking $F_{a}(x)=F_{\xi,a}(x)$ with $\xi=\frac{\Delta^{2}}{200}r$, we have the following bound (recall (2.6)):
@@
Moreover, since $\left\lvert\mathbf{x}_{D^{c}}\right\rvert=\left\lvert\mathbf{y}_{D^{c}}\right\rvert\geq 1$, we have $P_{I_{\leq q+\frac{1}{2}}}(H^{D,D^{c}})\bar{P}_{I_{\leq q-\frac{1}{2}}}(H^{D})=0$. Since
@@
$\hskip 20.0pt\leq\left\lvert\left(R^{D,D^{c}}_{z}P_{I_{\leq q-\frac{1}{2}}}(H^{D})\right)({\bf x},{\bf y})\right\rvert+\left\lvert\left(F_{a}(H^{D,D^{c}}-E)P_{I_{\leq q-\frac{1}{2}}}(H^{D})\right)({\bf x},{\bf y})\right\rvert$
@@
×$\hskip 199.16928pt\times\left\lvert\widehat{G}^{D}_{q-\frac{1}{2},\nu}(\mathbf{u}_{D},\mathbf{v}_{D})\right\rvert\left\lvert\pi^{D}_{\nu}(\mathbf{v}_{D},\mathbf{y}_{D})\right\rvert$
$\hskip 20.0pt\leq C_{q}\sum_{{\mathbf{u}_{D}}\in\mathcal{P}_{N_{D}}(D)}\sum_{{\mathbf{v}_{D}}\in\mathcal{P}_{{N_{D}},\lceil q-\frac{1}{2}\rceil}(D)}\mathrm{e}^{-cd_{H}^{\Lambda}\left({\bf x},\mathbf{u}_{D}\cup\mathbf{y}_{D^{c}}\right)}\mathrm{e}^{-c_{q-\frac{1}{2}}\left\lvert{\mathbf{u}_{D}}-{\mathbf{v}_{D}}\right\rvert_{1}}\left\lvert\pi^{D}_{\nu}(\mathbf{v}_{D},\mathbf{y}_{D})\right\rvert,$
where we used (3.16) and (C.3) for the last inequality and $N_{D}=\left\lvert\mathbf{x}_{D}\right\rvert$.
@@
$\hskip 20.0pt\leq C_{q}{\sum_{{\mathbf{u}_{D}}\in\mathcal{P}_{N_{D}}(D)}\sum_{{\mathbf{v}_{D}}\in\mathcal{P}_{{N_{D}},\lceil q-\frac{1}{2}\rceil}(D)}\mathrm{e}^{-cd_{H}^{\Lambda}{({\bf x},\mathbf{u}_{D}\cup\mathbf{y}_{D^{c}})}}\mathrm{e}^{-c_{q-\frac{1}{2}}\left\lvert{\mathbf{u}_{D}}-{\mathbf{v}_{D}}\right\rvert_{1}}}$
×$\hskip 170.71652pt\times\sum_{\nu\in\sigma_{q-\frac{1}{2}}(H^{D})}\left\lvert\pi^{D}_{\nu}(\mathbf{v}_{D},\mathbf{y}_{D})\right\rvert$
$\hskip 20.0pt\leq C_{q}{\sum_{{\mathbf{u}_{D}}\in\mathcal{P}_{N_{D}}(D)}\sum_{{\mathbf{v}_{D}}\in\mathcal{P}_{{N_{D}},\lceil q-\frac{1}{2}\rceil}(D)}\mathrm{e}^{-cd_{H}^{\Lambda}{({\bf x},\mathbf{u}_{D}\cup\mathbf{y}_{D^{c}})}}\mathrm{e}^{-c_{q-\frac{1}{2}}\left\lvert{\mathbf{u}_{D}}-{\mathbf{v}_{D}}\right\rvert_{1}}}\mathcal{Q}^{D}_{q-\frac{1}{2}}(\mathbf{v}_{D},\mathbf{y}_{D}).$
@@
$d_{H}^{\Lambda}({\bf x},\mathbf{u}_{D}\cup\mathbf{y}_{D^{c}})+\left\lvert{\mathbf{u}_{D}}-{\mathbf{v}_{D}}\right\rvert_{1}+d_{H}^{D}(\mathbf{v}_{D},\mathbf{y}_{D})\geq d_{H}^{\Lambda}(\mathbf{x},\mathbf{y}),$ (4.43)
@@
$d_{H}^{\Lambda}({\bf x},\mathbf{u}_{D}\cup\mathbf{y}_{D^{c}})+\left\lvert{\mathbf{u}_{D}}-{\mathbf{v}_{D}}\right\rvert_{1}+d_{H}^{D}(\mathbf{v}_{D}\cup\mathbf{y}_{D})\geq d_{H}^{\Lambda}({\bf x},\mathbf{u}_{D}\cup\mathbf{y}_{D^{c}})+d_{H}^{D}(\mathbf{u}_{D},\mathbf{y}_{D})$ (4.45)
$\quad=d_{H}^{\Lambda}({\bf x},\mathbf{u}_{D}\cup\mathbf{y}_{D^{c}})+d_{H}^{\Lambda}(\mathbf{u}_{D}\cup\mathbf{y}_{D^{c}},\mathbf{y})\geq d_{H}^{\Lambda}(\mathbf{x},\mathbf{y}),$

### v3

@@ 3. Basic features of the XXZ spin chain
$d_{1}^{D}(\mathbf{u},\mathbf{v})+d_{H}^{D}\left({\bf v},{\bf w}\right)+d_{1}^{D}({\bf w},{\bf y})\geq d_{H}^{D}({\bf x},{\bf y}),$ (4.14)
$d_{1}^{D}(\mathbf{u},\mathbf{v})+d_{H}^{D}\left({\bf v},{\bf w}\right)+d_{1}^{D}({\bf w},{\bf y})\geq d_{1}^{D}(\mathbf{u},\mathbf{v})+\frac{1}{2}\left(d_{H}^{D}\left({\bf v},{\bf y}\right)+d_{1}^{D}({\bf w},{\bf y})\right).$
@@
$\sum_{{\bf u}\in\mathcal{P}_{N}(D):x\in{\bf u}}\ \sum_{{\bf v},{\bf w}\in\mathcal{P}_{N,\lceil q\rceil}(D)}\mathrm{e}^{-c_{q}d_{1}^{D}(\mathbf{u},\mathbf{v})}\mathrm{e}^{-c_{q}d^{D}_{H}\left({\bf v},{\bf w}\right)}\mathrm{e}^{-c_{q}d_{1}^{D}({\bf w},{\bf y})}$ (4.15)
$\qquad\leq e^{-\frac{c_{q}}{2}d_{H}^{D}({\bf x},{\bf y})}\sum_{{\bf u}\in\mathcal{P}_{N}(D):}\ \sum_{{\bf v},{\bf w}\in\mathcal{P}_{N,\lceil q\rceil}(D)}e^{-\frac{c_{q}}{2}d_{1}^{D}(\mathbf{u},\mathbf{v})}e^{-\frac{c_{q}}{4}d_{H}^{D}\left({\bf v},{\bf w}\right)}e^{-\frac{c_{q}}{4}d_{1}^{D}({\bf w},{\bf y})}$
@@
using $\mathbf{y}\in\mathcal{P}_{N,\lceil q\rceil}(D)$, (B.10) and (B.3) in the last step.
@@
using a large deviation estimate. Let $\bar{\mu}=\mathbb{E}\left\{\omega_{0}\right\}$, and assume $N\lambda\bar{\mu}>2\lceil q\rceil\left(1-\tfrac{1}{\Delta}\right)$. Then the standard large deviations estimate (recall (3.2)) gives
@@
where we used $\left\lvert\mathcal{P}_{N,\lceil q\rceil}(S)\right\rvert=\operatorname{tr}Q_{\leq k}^{S}$, (3.16), and $\left\lvert S\right\rvert\leq N(2N+1)$. Moreover, on the complementary event $\left(\mathcal{E}_{N}^{S}\right)^{c}$ we have
@@
so we can use [15, Proposition 4.1] to obtain the Combes-Thomas bound
$\sup_{z\in\mathds{H}_{\lceil q\rceil}}\left\lvert G^{S}_{z}({\bf x},{\bf y})\right\rvert\leq C_{q}e^{-c_{q}d_{1}^{S}({\bf x},{\bf y})}\quad\text{for all}\quad\mathbf{x},\mathbf{y}\in\mathcal{P}_{N,\lceil q\rceil}(S).$ (4.22)
To show (4.17), we observe that for $\nu\in\sigma_{q}(H^{D})$ we have
$\pi_{{}_{\nu}}\phi_{\bf x}=\pi_{\nu}\left(H^{S,S^{c}}-\nu\right)R_{\nu}^{S,S^{c}}\phi_{\bf x}=\pi_{\nu}\Gamma^{S}R_{\nu}^{S,S^{c}}\phi_{\bf x},$ (4.23)
and the construction of $S$ yields
@@
It follows that on the complementary event $\left(\mathcal{E}_{N}^{S}\right)^{c}$ we have
$\sum_{\nu\in\sigma_{q}(H^{D})}\left\lvert\langle\phi_{\bf y},\pi_{\nu}\phi_{\bf x}\rangle\right\rvert$ $\leq\sum_{\nu\in\sigma_{q}(H^{D})}\sum_{{\bf u}\in\mathcal{P}_{N}^{\partial_{ex}}(D)}\left\lvert\langle\phi_{\bf y},\pi_{\nu}\phi_{\bf u}\rangle\right\rvert\left\lvert\langle\phi_{\bf u},\Gamma^{S}P_{+}^{S^{c}}R_{\nu}^{S}\phi_{\bf x}\rangle\right\rvert$ (4.25)
$\leq C_{q}\sum_{\nu\in\sigma_{q}(H^{D})}\sum_{{\bf u}\in\mathcal{P}_{N}^{\partial_{ex}}(D)}e^{-c_{q}d_{1}^{D}({\bf u},{\bf x})}\left\lvert\langle\phi_{\bf y},\pi_{\nu}\phi_{\bf u}\rangle\right\rvert$,
@@
$\mathcal{P}_{N}^{\partial_{ex}}(D)=\mathcal{P}_{N}(D,\partial_{ex}^{D}S)=\left\{{\bf u}\in\mathcal{P}_{N}(D),{\bf u}\cap\partial_{ex}^{D}S\neq\emptyset\right\}.$ (4.26)
@@
$\sum_{\nu\in\sigma_{q}(H^{D})}\left\lvert\langle\phi_{\bf y},\pi_{\nu}\phi_{\bf x}\rangle\right\rvert\leq C_{q}\sum_{{{\bf u}\in\mathcal{P}_{N}^{\partial_{ex}}(D)}}e^{-c_{q}d_{1}^{D}({\bf u},{\bf x})}.$ (4.28)
Since for any ${\bf u}\in\mathcal{P}_{N}^{\partial_{ex}}(D)$ we have $d_{1}^{D}({\bf u},{\bf x})\geq N$, we deduce from (4.28) and (B.3) (recall $\mathbf{y}\in\mathcal{P}_{N,\lceil q\rceil}(D)$), that on $\mathcal{E}_{N}^{c}$ we have
@@
$\pi_{\nu}^{D}(\mathbf{x},\mathbf{y})=C_{q}\sum_{\mathbf{u},\mathbf{v}\in\mathcal{P}_{N,\lceil q\rceil}(D)}\widehat{G}_{\lceil q\rceil,\nu}^{D}(\mathbf{x},\mathbf{u})\pi_{\nu}^{D}(\mathbf{u},\mathbf{v}){\widehat{G}}_{\lceil q\rceil,\nu}^{D}(\mathbf{v},\mathbf{y})$ (4.32)
@@
$\mathcal{Q}^{D}(\mathbf{x},\mathbf{y})$ $\leq\sum_{\mathbf{u},\mathbf{v}\in\mathcal{P}_{N,\lceil q\rceil}(D)}\left\lvert\widehat{G}^{D}_{\lceil q\rceil,\nu}(\mathbf{x},\mathbf{u})\right\rvert\mathcal{Q}^{D}(\mathbf{u},\mathbf{v})\left\lvert\widehat{G}^{D}_{\lceil q\rceil,\nu}(\mathbf{v},\mathbf{y})\right\rvert$ (4.33)
$\leq C_{q}\sum_{\mathbf{u},\mathbf{v}\in\mathcal{P}_{N,\lceil q\rceil}(D)}\mathrm{e}^{-c_{q}d_{1}^{D}({\bf x},{\bf u})}\mathcal{Q}^{D}(\mathbf{u},\mathbf{v})\mathrm{e}^{-c_{q}d_{1}^{D}({\bf v},{\bf y})},$
@@
$\mathbb{E}_{D}\left\{\mathcal{Q}^{D}(\mathbf{x},\mathbf{y}\right\}$ $\leq C_{q}N^{2\lceil q\rceil}\sum_{\mathbf{u},\mathbf{v}\in\mathcal{P}_{N,\lceil q\rceil}(D)}\mathrm{e}^{-c_{q}d_{1}^{D}({\bf x},{\bf u})}\mathrm{e}^{-c_{q}d_{H}^{D}(\mathbf{u},\mathbf{v})}\mathrm{e}^{-c_{q}d_{1}^{D}({\bf v},{\bf y})}$ (4.34)
$\leq C_{q}\mathrm{e}^{-c^{\prime}_{q}d_{H}^{D}(\mathbf{x},\mathbf{y})}\sum_{\mathbf{u},\mathbf{v}\in\mathcal{P}_{N,\lceil q\rceil}(D)}\mathrm{e}^{-c^{\prime}_{q}d_{1}^{D}({\bf x},{\bf u})}\mathrm{e}^{-c^{\prime}_{q}\mathrm{d}_{1}^{D}({\bf v},{\bf y})}$
@@
where we used (B.2) twice. This concludes the proof. ∎
@@
[…]
$\sup_{z\in\mathds{H}_{q}}\mathbb{E}_{\Lambda}\left\{\left\lvert G^{D,D^{c}}_{z}({\bf x},{\bf y})\right\rvert^{s}\right\}\leq C_{q}\left\lvert D\right\rvert^{2\lceil q\rceil}\mathrm{e}^{-c_{q-\frac{1}{2}}{d_{H}^{\Lambda}({\bf x},{\bf y})}}.$ (4.35)
@@
Let $\mathbf{x},\mathbf{y}\in\mathcal{P}_{N,\lceil q\rceil}(\Lambda)$, and consider a finite connected set $D\subset\Lambda$ satisfying $(\mathbf{x}\cup\mathbf{y})_{D}\neq\emptyset$ and $(\mathbf{x}\cup\mathbf{y})_{D^{c}}\neq\emptyset$. We only need to consider the case $1\leq\left\lvert\mathbf{x}_{D}\right\rvert=\left\lvert\mathbf{y}_{D}\right\rvert\leq N-1$, as otherwise $G^{D,D^{c}}_{z}({\bf x},{\bf y})=0$.
@@
Given $z\in\mathds{H}_{q}$, let $E=\operatorname{Re}z$ and $a=\operatorname{Im}z$. Setting $r=d_{H}^{\Lambda}\left({\bf x},{\bf y}\right)-1$, and taking $F_{a}(x)=F_{\xi,a}(x)$ with $\xi=\frac{\Delta^{2}}{50}r$, we have the following bound (recall (2.6)):
@@
Moreover, as $\left\lvert\mathbf{x}_{D^{c}}\right\rvert=\left\lvert\mathbf{y}_{D^{c}}\right\rvert\geq 1$, we have $P_{I_{\leq q+\frac{1}{2}}}(H^{D,D^{c}})\bar{P}_{I_{\leq q-\frac{1}{2}}}(H^{D})=0$. Since
@@
$\qquad\leq\left\lvert\left(R^{D,D^{c}}_{z}P_{I_{\leq q-\frac{1}{2}}}(H^{D})\right)({\bf x},{\bf y})\right\rvert+\left\lvert\left(F_{a}(H^{D,D^{c}}-E)P_{I_{\leq q-\frac{1}{2}}}(H^{D})\right)({\bf x},{\bf y})\right\rvert$
@@
$\hskip 199.16928pt\times\left\lvert\widehat{G}^{D}_{q-\frac{1}{2},\nu}(\mathbf{u}_{D},\mathbf{v}_{D})\right\rvert\left\lvert\pi^{D}_{\nu}(\mathbf{v}_{D},\mathbf{y}_{D})\right\rvert$
$\hskip 20.0pt\leq C_{q}\sum_{{\mathbf{u}_{D}}\in\mathcal{P}_{N_{D}}(D)}\sum_{{\mathbf{v}_{D}}\in\mathcal{P}_{{N_{D}},\lceil q-\frac{1}{2}\rceil}(D)}\mathrm{e}^{-cd_{H}^{\Lambda}\left({\bf x},\mathbf{u}_{D}\cup\mathbf{y}_{D^{c}}\right)}\mathrm{e}^{-c_{q-\frac{1}{2}}d_{1}^{D}({\mathbf{u}_{D}},{\mathbf{v}_{D}})}\left\lvert\pi^{D}_{\nu}(\mathbf{v}_{D},\mathbf{y}_{D})\right\rvert$,
where $N_{D}=\left\lvert\mathbf{x}_{D}\right\rvert$ and we used (3.24) and (C.3) for the last inequality.
@@
$\hskip 20.0pt\leq C_{q}{\sum_{{\mathbf{u}_{D}}\in\mathcal{P}_{N_{D}}(D)}\sum_{{\mathbf{v}_{D}}\in\mathcal{P}_{{N_{D}},\lceil q-\frac{1}{2}\rceil}(D)}\mathrm{e}^{-cd_{H}^{\Lambda}{({\bf x},\mathbf{u}_{D}\cup\mathbf{y}_{D^{c}})}}\mathrm{e}^{-c_{q-\frac{1}{2}}d_{1}^{D}({\mathbf{u}_{D}},{\mathbf{v}_{D}})}}$
$\hskip 170.71652pt\times\sum_{\nu\in\sigma_{q-\frac{1}{2}}(H^{D})}\left\lvert\pi^{D}_{\nu}(\mathbf{v}_{D},\mathbf{y}_{D})\right\rvert$
$\hskip 20.0pt\leq C_{q}{\sum_{{\mathbf{u}_{D}}\in\mathcal{P}_{N_{D}}(D)}\sum_{{\mathbf{v}_{D}}\in\mathcal{P}_{{N_{D}},\lceil q-\frac{1}{2}\rceil}(D)}\mathrm{e}^{-cd_{H}^{\Lambda}{({\bf x},\mathbf{u}_{D}\cup\mathbf{y}_{D^{c}})}}\mathrm{e}^{-c_{q-\frac{1}{2}}d_{1}^{D}({\mathbf{u}_{D}},{\mathbf{v}_{D}})}}\mathcal{Q}^{D}_{q-\frac{1}{2}}(\mathbf{v}_{D},\mathbf{y}_{D}).$
@@
$d_{H}^{\Lambda}({\bf x},\mathbf{u}_{D}\cup\mathbf{y}_{D^{c}})+d_{1}^{D}({\mathbf{u}_{D}},{\mathbf{v}_{D}})+d_{H}^{D}(\mathbf{v}_{D},\mathbf{y}_{D})\geq d_{H}^{\Lambda}(\mathbf{x},\mathbf{y}),$ (4.43)
@@
$d_{H}^{\Lambda}({\bf x},\mathbf{u}_{D}\cup\mathbf{y}_{D^{c}})+d_{1}^{D}({\mathbf{u}_{D}},{\mathbf{v}_{D}})+d_{H}^{D}(\mathbf{v}_{D}\cup\mathbf{y}_{D})\geq d_{H}^{\Lambda}({\bf x},\mathbf{u}_{D}\cup\mathbf{y}_{D^{c}})+d_{H}^{D}(\mathbf{u}_{D},\mathbf{y}_{D})$ (4.45)
$\quad=d_{H}^{\Lambda}({\bf x},\mathbf{u}_{D}\cup\mathbf{y}_{D^{c}})+d_{H}^{\Lambda}(\mathbf{u}_{D}\cup\mathbf{y}_{D^{c}},\mathbf{y})\geq d_{H}^{\Lambda}(\mathbf{x},\mathbf{y}).$